Math 122  University of South Carolina
D. Meade  Fall 1997

Math 122 (Section 2) – Business Calculus

Instructor Professor Doug Meade

Office Hours: MWF 10:00–11:00 and by prior appointment
Office:  LeConte College 300E
Phone:  777-6183

 E-mail:  meade@math.sc.edu

WWW URL  http://www.math.sc.edu/~meade/math122-F97/

Meeting Times MWF 9:05am– 9:55am, CLS 104

Text Brief Calculus, preliminary edition, by Hughes-Hallet, Gleason, et al., John Wiley & Sons, 1997.

Prerequisite Qualification through placement, or a grade of C or better in Math 111 or 115.

Overview One of the main objectives of this course is for you to understand the basic concepts of calculus well enough to know when, how, and why to apply them in real-world situations and to be able to interpret and communicate the results.

To achieve this goal will require practice at a variety of numerical, graphical, and analytic methods. In particular, being proficient only with the mechanical manipulations is not sufficient for are successful completion of this course.

You will also be expected to develop and practice your verbal and written communication skills. You will be required to work in groups during class and will have ample opportunities to practice these skills. It is hoped that you will also find a group to work with outside of class.

The preface of the book (particularly page xi) is particularly well-written. (Pages vii – xi provide additional detail and insight into the methods you will encounter in this course.) Read the preface!

Use of Technology This course will make regular use of a graphing calculator. The TI-83 is the preferred calculator. A TI-82 is similar enough that it should be sufficient for our needs. (If you choose to use a different calculator, I do not promise to be able to assist you.) One of the objectives of this course is for you to become proficient in the appropriate use of your calculator. This includes recognizing when the calculator should be used and when hand manipulations are more efficient.

Course Content We will be discussing the first six chapters of the text. This will bring you into contact with the standard topics of calculus. That is, we will be looking at the way quantities change (differentiation) and accumulate (integration).

Study Hints Before each class, you should both review the material from recent sections and read the section to be discussed that day. This will allow you to both understand my presentation of new material and identify questions that you need to resolve within your in-class group time.

Grading Your grade in this course will be based on your performance on quizzes, three (3) mid-term exams, and a final exam. The weights assigned to each of these components will be:

Quizzes  10%
Mid-term exams (3) 60%
Final exam  30%

Course grades will be determined according to the following scale:

A 90 –100
B 80 – 89
C 70 – 79
D 60 – 69
F 0 – 59

Note that the deadline to drop this course with a grade of W is \underline{Thursday, October 2, 1997}.

Exams There will be three (3) exams during the semester. Tentative dates and topics for these exams are:

Wednesday, September 24  Chapters 1 and 2
Wednesday, October 29  Chapters 3 and 4
Monday, November 24  Chapters 5 and 6

There will be no make-up exams. If you miss one exam due to a documented reason of illness, family emergency or participation in a University sponsored event, your score on the final exam will be used to replace the missing exam score. Excuses such as oversleeping, forgetting the time or location of the exam, and lack of studying are explicitly noted as unacceptable grounds for missing an exam.

A comprehensive final will be given at \underline{9:00a.m.} on \underline{Wednesday, December 10, 1997}.

Quizzes On average, there will be one or two quizzes each week. The quiz questions will be based on the homework problems but generally shorter. Each quiz will be graded on a five (5) point scale. Your quiz grade will be based on the sum of your twelve (12) highest quiz scores. There are no makeup quizzes.

Homework A minimum set of homework problems will be announced for each section that we discuss. You will be given some time to work on these problems during class, but no homework will be collected. You are responsible for working and understanding all of the assigned problems.

Attendance Regular class attendance is important. Consistent with the USC Undergraduate Bulletin, a grade penalty may be applied to any student missing more than four classes (10%) during the semester.

Academic Honesty Cheating and plagiarism will not be tolerated in this course. You will be working in groups in class and are encouraged to discuss homework problems with others. You will have to take all quizzes and exams on your own. Violations of this policy will be dealt with in a manner consistent with University guidelines.